Circular No. 8955

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2467 CYGNI D. K. Lynch, R. W. Russell, and R. J. Rudy, The Aerospace Corpora-

tion; and C. E. Woodward, University of Minnesota, report 0.8- to 5.5-μm spectroscopy of V2467 Cyg (cf. IAUC 8905) on June 15 UT using SpeX on the Infrared Telescope Facility. The nova has persisted in a very high state of excitation, and its spectrum is dominated by coronal lines. The [Mg VIII] line at 3.03 μm is the brightest feature in the infrared spectrum, and lines from four separate ionization states of silicon are present ([Si VI], [Si VII], [Si IX], [Si X]). Other species displaying coronal lines include [P VIII], [S VIII], [S IX], [Ca VII], [Ca VIII], and possibly [Ni X]. The thermal infrared continuum remains free of dust emission.

COMETS C/2008 D5–D9 (SOHO) Further to IAUC 8953, additional apparent comets have been found

on SOHO website images; all are Kreutz sungrazers except for C/2008 D6 (Meyer group). C/2008 D5 and C/2008 D7 were small, stellar in appearance, and of mag ∼ 7. C/2008 D6 was elongated but not diffuse (mag ∼ 7). C/2008 D8 was slightly diffuse and very faint (mag 7.5). C/2008 D9 was slightly diffuse and small (mag 7) in C3 images, but very diffuse and very faint in C2 images. An asterisk in the instrumentation column below indicates that astrometry was also obtained from STEREO-A/SECCHI HI1 images (see IAUC 8926).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 D5 Feb. 19.029 22^{h}23.^{m}1 −12^{o}41′ C3  HS 2008-L19
C/2008 D6  19.171 22 12.8 −10 00 C2  TH 2008-L19
C/2008 D7  25.554 22 51.5 − 9 54 C3*  BZ 2008-L19
C/2008 D8  24.554 22 43.7 −10 03 C3*  BZ 2008-L26
C/2008 D9  29.238 22 08.4 − 9 00 C3/2* BZ 2008-L26

COMETS C/2008 E5 AND C/2008 E6 (STEREO) Further to IAUC 8926, additional very faint (mag ∼ 11) Kreutz sun-

grazers have been found on STEREO website images:

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 E5 Mar. 5.517 0^{h}54.^{m}2 + 1^{o}44′ HI  AW 2008-K29
C/2008 E6  15.017 1 28.6 + 5 05 HI  RK 2008-K29
2008 June 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green